OpenStudy (anonymous):

Solve the triangle ABC in which α=48°32' , a=38.87 cm & c=24.93 cm . . . ANSWER= b=50.59, β=102°45' , γ=28°43'

OpenStudy (anonymous):

????

hartnn (hartnn):

sine rule will directly give you γ

OpenStudy (anonymous):

\[\frac{ a }{ \sin \alpha }=\frac{ c }{ \sin \gamma }\]

OpenStudy (anonymous):

what will be the formula of sin gamma

hartnn (hartnn):

then cos rule will give you b c^2 = a^2+b^2 -2ab cos γ then sine rule again will give you β. sin gamma is only unknown there.... you'll get sin gamma =.... gamma = sin^-1 (....)

OpenStudy (anonymous):

what will be the formula of sin gamma??

OpenStudy (anonymous):

sin gamma = sin alpha x c /a is this the formula???

hartnn (hartnn):

oh, that...yes (c/a) * sin aplha

OpenStudy (anonymous):

its coming wrong

OpenStudy (anonymous):

my calculator is giving a math error

hartnn (hartnn):

....sin gamma is coming > 1 ??

OpenStudy (anonymous):

it is coming < 1 0_0

hartnn (hartnn):

sin inverse of 0.479

hartnn (hartnn):

28.621 degrees

OpenStudy (anonymous):

I GOT IT I WAS TYPING WRONG

hartnn (hartnn):

okk..so got b and beta also ?

OpenStudy (anonymous):

we can also find beta by alpha + beta + gamma = 180 degrees beta = 180 - (alpha +gamma)

hartnn (hartnn):

good!

hartnn (hartnn):

i mean better :)